# https://vl2g.github.io/projects/IndicPhotoOCR/

### v1

@@ Abstract
To address these gaps, we introduce the Bharat Scene Text Dataset (BSTD) – a large-scale and comprehensive benchmark for studying Indian Language Scene Text Recognition. It comprises more than 100K words that span 11 Indian languages and English, sourced from over 6,500 scene images captured across various linguistic regions of India. The dataset is meticulously annotated and supports multiple scene text tasks, including: (i) Scene Text Detection, (ii) Script Identification, (iii) Cropped Word Recognition, and (iv) End-to-End Scene Text Recognition. We evaluated state-of-the-art models originally developed for English by adapting (fine-tuning) them for Indian languages. Our results highlight the challenges and opportunities in Indian language scene text recognition. We believe that this dataset represents a significant step toward advancing research in this domain. All our models and data are open source.
@@
Figure 1: Scene Text Understanding: Case of India. This figure showcases typical street scenes from northern, eastern, western, and southern parts of India, highlighting the country’s vast linguistic diversity. With 11 languages (five of them: Hindi, English, Bengali Gujarati and Tamil are shown here) including English commonly featured on signboards, scene text understanding in India presents unique challenges. While significant progress has been made in English Scene Text Recognition, open-source comprehensive effort for Indian language scene text understanding, including large public datasets and comprehensive models, are still limited. Our work seeks to address this gap and advance the field.
@@ 1 Introduction
Scene Text Recognition, especially for English, has advanced significantly in the last decade, thanks to a series of efforts on dataset creation [1, 2, 3, 4, 5, 6, 7, 8, 9, 10] and model development [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. In contrast, scene text recognition for Indian languages has been extremely underexplored and possibly more challenging (refer to Fig. 1). Note that Indian languages are used by more than 1.4 billion people, which is approximately 18% of the global population. Despite its usage and popularity, we observe that scene text recognition in Indian languages lacks a comprehensive benchmark and open-source models, except for some early efforts [26, 27, 28, 29, 30, 31, 32, 33] that laid important groundwork, but did not aim for a comprehensive coverage of tasks or languages. We aim to fill this gap through our work.
To accelerate advancements in Indian Language Scene Text Recognition, we present a comprehensive benchmark called the Bharat Scene Text Dataset (BSTD) [^1] India, traditionally known as Bharat, reflects the nation’s rich cultural and linguistic diversity, which inspired the naming of our dataset.. The BSTD comprises 6.5K scene images that feature more than 100K words in 11 Indian languages, along with English. These images, sourced from Wikimedia [^2] https://www.wikimedia.org/, span various signboards and accidental background text from different Indian cities and towns, ensuring that the dataset captures the linguistic diversity of India. We obtain a polygon-level bounding box for each visible text in the image and the text it contains using rigorous manual annotation. The dataset is publicly available for download here [^3] https://github.com/Bhashini-IITJ/BharatSceneTextDataset.
The BSTD can be used for the following scene text tasks: (i) scene text detection, (ii) script identification, (iii) cropped word recognition, and (iv) end-to-end scene text recognition. In this work, we benchmark the BSTD using our new pipeline, [Uncaptioned image], which integrates state-of-the-art architectures across its modules. We firmly believe that these benchmarks will help accelerate advancements in Indian language scene text recognition.
@@
ICDAR 2003 [1] 509 2,268 1
SVT [2] 350 725 1
IIIT-5K [3] - 5,000 1
ICDAR 2013 [5] 462 5,003 1
ICDAR 2015 [6] 1,500 6,545 1
CUTE80 [7] 80 288 1
Total-Text [8] 1,555 11,459 1
IIIT-TL-STR [30] 440 3,035 2
IIIT-ILST [27] - 3,168 3
MLT-17 [34] 18,000 96,000 9
MLT-19 [10] 20,000 191,000 10
IIIT-indicSTR12∗∗ [31] - 27,000 12
BSTD (This Work) 6,582 1,06,478 12
Table 1: Comparison of the Bharat Scene Text Dataset (BSTD) with existing scene text datasets. BSTD supports all three core tasks—text detection, script identification, and recognition (and, thereby end-to-end evaluation) across 12 Indian languages, including English. It is the only publicly available dataset for Indian languages that enables comprehensive end-to-end scene text understanding. : MLT-17 and MLT-19 are not focused on Indian languages, they only include Hindi and Bengali. **: IIIT-IndicSTR12 is primarily annotated for cropped word recognition and can not be studied to evaluate end-to-end scene text recognition as ours.
@@
    We introduce the Bharat Scene Text Dataset (BSTD) – a large-scale benchmark for Indian language scene text recognition, consisting of 6,582 scene images featuring 1,06,478 words in 11 Indian languages and English. Each image is manually annotated with polygon-level bounding boxes and corresponding text and script, ensuring high-quality data for research and applications. (Section 2)
@@
    We evaluated BSTD using state-of-the-art techniques, including off-the-shelf methods and fine-tuned models on synthetic and/or BSTD train data (as applicable). Detailed benchmark results are provided for all key tasks, i.e., text detection, script identification, cropped word recognition, and end-to-end scene text recognition. These evaluations establish a strong baseline for future research and advancements in multilingual scene text recognition, particularly for Indian languages. (Section 4)
@@
    We introduce an open source toolkit, namely [Uncaptioned image], designed to detect, identify, and recognize text in English and 11 Indian languages. The toolkit is easy to install and replicate, making it accessible to researchers and developers. (Section 5)
@@ 2 Bharat Scene Text Dataset
India, often referred to as Bharat (the Indian name for the country), represents a rich and diverse cultural and linguistic landscape, as illustrated in Fig. 2. In the context of scene text understanding, the vast multilingual environment poses unique challenges, as the scripts, writing styles, and text formats vary significantly between different regions. In addition, the identification of scripts becomes extremely important. Although there has been some effort in collecting and benchmarking Indian language scene text understanding tasks, it falls short compared to efforts for scene text understanding in English [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. Please refer to Table 1 for a detailed comparison of the Bharat Scene Text Dataset compared to other related scene text datasets.
To address this gap, we present the Bharat Scene Text Dataset, a novel and comprehensive dataset specifically curated to represent the linguistic diversity of India. By using the term “Bharat”, we not only acknowledge the nation’s deep-rooted cultural identity but also highlight our focus on supporting scene text understanding across 11 major Indian languages. This dataset contains 6,582 scene images and 1,06,478 word images in total. This dataset and associated annotations can be downloaded from our project website [^4] https://vl2g.github.io/projects/IndicPhotoOCR/. In this section, we give details of dataset curation and annotation, supported tasks, train-test splits followed by dataset analysis.
@@ 2.1 Dataset Curation and Annotation
Figure 3: Pipeline for constructing a multilingual scene-text dataset from Wikimedia Commons images. The process begins with query construction using combinations of Indian city names and place-related keywords to retrieve relevant images via the Wikimedia Commons API. A two-stage filtering process using DBnet [19] (pre-trained on Indic Synthetic Data) removes non-textual images and manually filters relevant ones. Scene-text detection annotations are generated using DBnet and refined through manual corrections. Finally, cropped word images are recognized using the PARSeq [22] model (trained on 11 Indic languages using synthetic data), with recognition errors corrected and script language tags added to produce the final annotated dataset.
@@ 2.1.1 Stage-1: Image collection
We compile a comprehensive list of cities and towns in various regions of India where 11 selected languages are spoken predominantly. Furthermore, we curate a catalog of public places, such as ‘bus stations’, ‘banks’, ‘schools’, ‘railway stations’, ‘hospitals’, ‘ATMs’, ‘metro stations’, ‘street signs’, ‘airport’, ‘shop entrance’, ‘billboards’, ’press conference’, etc. By combining the city and place names as search queries, e.g. ‘Delhi Railway Station’, we retrieved images along with their metadata from the Wikimedia Commons API [^5] https://commons.wikimedia.org/w/api.php. It should be noted that not all retrieved images contain scene text at this stage, which requires subsequent filtering. At this stage, we obtain 50K images.
@@ 2.1.2 Stage-2: Image filtering
We filter the images harvested in the previous step using the pre-trained scene text detection method [19]. To this end, we run this detector and obtain text even with low confidence. We retain images containing at least one word, while others were discarded. In addition, we ask human annotators to review the filtered set to eliminate false positives. The final data set comprises 6,582 images that contain scene text in 11 Indian languages and English.
@@ 2.1.3 Stage-3: Semi-automatic annotation
We annotate the final filtered images for scene text detection in two steps to reduce human annotation efforts. We obtain bounding box annotation using a fine-tuned DBnet model on synthetic Indian language data [^6] Refer to Section-3.3.1 for details of synthetic data generation. Since these automatically obtained bounding boxes are not always perfect, they are refined by our hired annotators using an in-house tool to edit, draw, or delete polygons around the scene text words.
Once the bounding boxes were obtained, they were assigned to the annotators and assigned to label each box with the script it contains. These annotators were well-versed in script identification and were also informed about the geographical origin of each image. With this context, they were able to accurately label each bounding box with the corresponding script. We then use PARSeq [22], fine-tuned on synthetic data for Indian languages, to read the text within each bounding box. The bounding boxes, along with the recognized text, are then provided to language expert annotators. These annotators review the text and correct any recognition errors or retain the text as is if it is accurate.
@@ 2.2 Annotator Details
We employed two distinct groups of annotators for this work. Initially, three human annotators for image filtering (Section 2.1.2) and bounding box annotation (Section 2.1.3). For the next two steps, text correction and script assignment (refer 2.1.3), we hired eleven language-specific experts, one for each target language. The annotators were proficient in reading, understanding, and writing their native language, as well as English. They were shown cropped words and asked to correct the pseudo-recognition annotations and assign the script. All final annotations underwent an additional quality check to eliminate potential errors and noise. The annotators were compensated at rates aligned with standard wages for a tier-2 city in India. The entire process–pseudo-annotation, manual correction, and quality verification took 10 months. Fig 4 shows the fine grained nature of the annotations done in this work.
@@ 2.3 Supported Tasks and Evaluation Protocols
    Scene Text Detection: where given an image, the task is to detect all the text instances in the image along with their bounding boxes. We provide polygon-level tight bounding boxes for each word image in the dataset. The popular metrics such as Precision, Recall and F1-score using TedEval evaluation [35] can be used to assess the performance of the method.
@@
    End-to-End Scene Text Recognition: This task is the closest to the natural setting. Here, given a scene image, the task is to recognize all the instances of text. Solving end-to-end Scene Text Recognition for Indian languages, often requires a pipeline of scene text detection, script identification followed by cropped word recognition. We use mean word recognition rate (WRR) and character recognition rate (CRR) over all the test images in the dataset. The WRR and CRR for an image are defined as follows:
@@
    where $S$, $D$, and $I$ with subscript $w$ and $c$ denote the number of substitutions, deletions, and insertions, respectively, at the word or character level. Further, $N_{w}^{\text{total}}$ and $N_{c}^{\text{total}}$ denote the total number of words and characters present in the ground-truth transcription in the image, respectively. It should be noted that due to false alarms (i.e., excessive insertions of words or characters in the recognized text) or in some rare cases of any missing annotation, the computed WRR or CRR values can become negative for few images. To keep the interpretation simple, we truncate such values to 0. Furthermore, standard WRR and CRR computations assume a fixed reading order; however, since reading order in scene text images can be subjective, we relax this constraint in our evaluation. Further, to make interpretation of results more clean, we also use standard precision, recall, and F1-score to evaluate end-to-end performance. Note that precision is defined as the number of correctly recognized words over the total number of recognized words, while recall is defined as the number of correctly recognized words over the total number of words in the ground truth.
@@
Table 2: Summary of image and detection annotation counts in the BSTD dataset, detailing the total, training, and test set distributions
@@
Table 3: Cropped word image data split for the BSTD dataset, showing the number of annotated instances, train images, and test images across multiple languages. Here, Others correspond to those words which are not from the listed languages here (e.g., Meitei and Urdu). We plan to include their annotation in subsequent versions of our dataset.
Figure 6: Distribution of the bounding box size in our dataset, in terms of pixels.

### v2

@@ Abstract
To address these gaps, we introduce the Bharat Scene Text Dataset (BSTD) – a large-scale and comprehensive benchmark for studying Indian Language Scene Text Recognition. It comprises more than 100K words that span 11 Indian languages and English, sourced from over 6,582 scene images captured across various linguistic regions of India. The dataset is meticulously annotated and supports multiple scene text tasks, including: (i) Scene Text Detection, (ii) Script Identification, (iii) Cropped Word Recognition, and (iv) End-to-End Scene Text Recognition. We evaluated state-of-the-art models originally developed for English by adapting (fine-tuning) them for Indian languages. Our results highlight the challenges and opportunities in Indian language scene text recognition. We believe that our dataset represents a significant step toward advancing research in this domain. All our models and data are open source.
@@
Figure 1: Scene Text Understanding: Case of India. This figure showcases typical street scenes from northern, eastern, western, and southern parts of India, highlighting the country’s vast linguistic diversity. With 11 languages including English commonly featured on signboards, scene text understanding in India presents unique challenges. Scene Text from five languages, namely Hindi, English, Bengali Gujarati and Tamil are shown here. While significant progress has been made in English Scene Text Recognition, open-source comprehensive effort for Indian language scene text understanding, including large public datasets and open-source models, are still limited. Our work seeks to address this gap toward advancing the field.
@@ 1 Introduction
Scene Text Recognition, especially for English, has advanced significantly in the last decade, thanks to a series of efforts on dataset creation [ICDAR2003, SVT, mishra2012scene, MSRA-TD500, ICDAR2013, ICDAR2015, CUTE80, Total-Text, SVT-P, MLT-19] and model development [cnn_detector_2014, CRNN2016, jaderberg2016reading, east, textspotter2018, craft, ASTER2019, ABCNet2020, dbnet, visionLAN2021, ABINet2021, parseq, ABINet++2023, textbpn++, clip4str]. In contrast, scene text recognition for Indian languages has been extremely underexplored and possibly more challenging (refer to Fig. 1). Note that Indian languages are used by more than 1.4 billion people, which is approximately 18% of the global population. Despite its usage and popularity, we observe that scene text recognition in Indian languages lacks a comprehensive benchmark and open-source models, except for some early efforts [recognition-ISI-umapada2010, IIIT_ILST, detection-ISI-Umapada2017, kakwani-etal-2020-indicnlpsuite, IIIT-TL-STR, IIIT-indicSTR12, VijayanCDK25] that laid important groundwork, but did not aim for a comprehensive coverage of tasks or languages. We aim to fill this gap through our work.
To accelerate advancements in Indian Language Scene Text Recognition, we present a comprehensive benchmark called the Bharat Scene Text Dataset (BSTD) [^1] India, traditionally known as Bharat, reflects the nation’s rich cultural and linguistic diversity, which inspired the naming of our dataset.. The BSTD comprises 6.5K scene images that feature more than 100K words in 11 Indian languages, along with English. These images that are sourced from Wikimedia [^2] https://www.wikimedia.org/, span various signboards and accidental background text from different Indian cities and towns, ensuring that the dataset captures the linguistic diversity of India. For each visible text instance in an image, we manually annotate a polygonal bounding region and provide its corresponding transcription. The dataset is publicly available for download [^3] https://github.com/Bhashini-IITJ/BharatSceneTextDataset.
The BSTD can be used for the following scene text tasks: (i) scene text detection, (ii) script identification, (iii) cropped word recognition, and (iv) end-to-end scene text recognition. In this work, we benchmark the BSTD using our new pipeline, [Uncaptioned image], which integrates state-of-the-art architectures in its modules. We firmly believe that these benchmarks will help accelerate advancements in Indian language scene text recognition.
Table 1: Comparison of the Bharat Scene Text Dataset (BSTD) with existing scene text datasets. BSTD supports all three core tasks - text detection, script identification, and recognition (and, thereby end-to-end evaluation) across 11 Indian languages along with English. It is the only publicly available dataset for Indian languages that enables comprehensive end-to-end scene text understanding. : MLT-17 and MLT-19 are not focused on Indian languages, they only include Hindi and Bengali. \$: IIIT-IndicSTR12 and #: IIIT-STR [GunnaSJ22] are primarily annotated for cropped word recognition; consequently, neither can be utilized to evaluate end-to-end scene text recognition as ours.
@@
ICDAR 2003 [ICDAR2003] 509 2,268 1
SVT [SVT] 350 725 1
IIIT-5K [mishra2012scene] - 5,000 1
ICDAR 2013 [ICDAR2013] 462 5,003 1
ICDAR 2015 [ICDAR2015] 1,500 6,545 1
CUTE80 [CUTE80] 80 288 1
Total-Text [Total-Text] 1,555 11,459 1
IIIT-TL-STR [IIIT-TL-STR] 440 3,035 2
IIIT-ILST [IIIT_ILST] - 3,168 3
MLT-17 [MLT-17] 18,000 96,000 9
MLT-19 [MLT-19] 20,000 191,000 10
IIIT-indicSTR12\$ [IIIT-indicSTR12] - 27,000 12
IIIT-STR# [GunnaSJ22] - 12,057 6
BSTD (This Work) 6,582 1,26,292 12
@@
    We introduce the Bharat Scene Text Dataset (BSTD) – a large-scale benchmark for Indian language scene text recognition, consisting of 6,582 scene images featuring 1,26,292 words in 11 Indian languages and English. Each image is manually annotated with polygon-level bounding boxes and corresponding transcription and script, ensuring high-quality data for research and applications. (Section 2)
@@
    We present a strong baseline and release corresponding open-source toolkit, namely [Uncaptioned image], designed to detect, identify, and recognize text in English and 11 Indian languages. The released toolkit is easy to install and replicate, making it accessible to researchers and developers. (Section 3 and Appendix-A)
@@
    We evaluated BSTD using state-of-the-art techniques, including off-the-shelf methods and fine-tuned models on synthetic and/or BSTD train data (as applicable). Detailed benchmark results including failure case analysis are provided for all key tasks, i.e., text detection, script identification, cropped word recognition, and end-to-end scene text recognition. These evaluations establish a strong baseline for future research and advancements in multilingual scene text recognition, particularly for Indian languages. (Section 4)
Figure 3: Pipeline for constructing a multilingual scene-text dataset from Wikimedia Commons images. The process begins with query construction using combinations of Indian city names and place-related keywords to retrieve relevant images via the Wikimedia Commons API. A two-stage filtering process using DBnet [dbnet] (pre-trained on Indic Synthetic Data) removes non-textual images and manually filters relevant ones. Scene-text detection annotations are generated using DBnet and refined through manual corrections. Finally, cropped word images are recognized using the PARSeq [parseq] model (trained on 11 Indic languages using synthetic data), with recognition errors corrected and script language tags added to produce the final annotated dataset.
@@ 2 Bharat Scene Text Dataset
India, often referred to as Bharat (the Indian name for the country), represents a rich and diverse cultural and linguistic landscape, as illustrated in Fig. 2. In the context of scene text understanding, the vast multilingual environment poses unique challenges, as the scripts, writing styles, and text formats vary significantly between different regions. In addition, the identification of scripts becomes extremely important. Although there has been some effort in collecting and benchmarking Indian language scene text understanding, such as [IIIT_ILST, IIIT-TL-STR, IIIT-indicSTR12, GunnaSJ22], it falls short compared to efforts for scene text understanding in English [ICDAR2003, SVT, mishra2012scene, MSRA-TD500, ICDAR2013, ICDAR2015, CUTE80, Total-Text, SVT-P, MLT-19]. To address this gap, we present the Bharat Scene Text Dataset, a novel and comprehensive dataset specifically curated to represent the linguistic diversity of India. By using the term “Bharat”, we not only acknowledge the nation’s deep-rooted cultural identity but also highlight our focus on supporting scene text understanding across 11 major Indian languages. This dataset contains 6,582 scene images and 1,26,292 word images in total. This dataset and associated annotations can be downloaded from our project website [^4] https://vl2g.github.io/projects/IndicPhotoOCR/. Please refer to Table 1 for a detailed comparison of the Bharat Scene Text Dataset compared to other related scene text datasets.
In the next section, we give details of dataset curation and annotation, supported tasks, train-test splits followed by dataset analysis.
@@ 2.1.1 Stage-1: Image collection
We compile a comprehensive list of cities and towns across various regions of India where 11 selected languages are spoken predominantly. Furthermore, we curate a catalog of public places, such as ‘bus stations’, ‘banks’, ‘schools’, ‘railway stations’, ‘hospitals’, ‘ATMs’, ‘metro stations’, ‘street signs’, ‘airport’, ‘shop entrance’, ‘billboards’, ’press conference’, etc. By combining the city and place names as search queries, e.g. ‘Delhi Railway Station’, ‘Kolkata Bookshop’, we retrieved images along with their metadata from the Wikimedia Commons API [^5] https://commons.wikimedia.org/w/api.php. It should be noted that not all retrieved images contain scene text, necessitating a subsequent filtering process. At this stage, we obtain $\approx$ 50K images.
@@ 2.1.2 Stage-2: Image filtering
We filter the images harvested in the previous step using the pre-trained scene text detection method [dbnet]. To this end, we run this detector and obtain text even with low confidence. We retain images containing at least one word, while others were discarded. Nevertheless, 53 images in the final dataset were found to contain no annotatable scene text. We deliberately retained this small fraction of non-text images to preserve the natural variability of real-world data collection pipelines, where some captured scenes may not ultimately contain visible text. The final data set comprises 6,582 images that contain scene text in 11 Indian languages and English.
@@ 2.1.3 Stage-3: Semi-automatic annotation
We annotate the final filtered images for scene text detection in two steps to reduce human annotation efforts. We obtain bounding box annotation using a fine-tuned DBnet model on synthetic Indian language data [^6] Refer to Section-3.3.1 for details of synthetic data generation. Since these automatically obtained bounding boxes are not always perfect, they were refined by our hired annotators using an internal tool to edit, draw, or delete polygons around the scene text words.
Once the bounding boxes were obtained, they were assigned to the annotators and tasked to label each box with the script it contains. These annotators were well-versed in script identification and were also informed about the geographical origin of each image. With this context, they were able to accurately label each bounding box with theier corresponding script. We then use PARSeq [parseq], fine-tuned on synthetic data for Indian languages, to read the text within each bounding box. The bounding boxes, along with the recognized text, are then provided to language expert annotators. These annotators review the text and correct any recognition errors or retain the text as is if it is accurate. In addition, for every image, we have included annotation on the lighting (captured during the day or night) and environment (captured outdoor or indoor) conditions.
Fig. 4 shows the fine grained nature of the annotations done in this work.
@@ 2.2 Annotator Details
We employed two distinct groups of annotators for this work. Initially, three human annotators for image filtering (Section 2.1.2) and bounding box annotation (Section 2.1.3). For the next two steps, text correction and script assignment (refer 2.1.3), we hired eleven language-specific experts, one for each target language. The annotators were proficient in reading, understanding, and writing their native language, as well as English. They were shown cropped words and asked to correct the pseudo-recognition annotations and assign the script. All final annotations underwent an additional quality check to eliminate potential errors and noise. The annotators were compensated at rates aligned with standard wages for a tier-2 city in India. The entire process–pseudo-annotation, manual correction, and quality verification took 10 months.
@@ 2.3 Supported Tasks and Evaluation Protocols
    Scene Text Detection: where given an image, the task is to detect all the text instances in the image along with their bounding boxes. We provide polygon-level tight bounding boxes for each word image in the dataset. The popular metrics such as Precision, Recall and F1-score using TedEval evaluation [tedeval] can be used to assess the performance of the scene text detectors.
@@
    End-to-End Scene Text Recognition: This task is the closest to the natural setting. Here, given a scene image, the task is to recognize all instances of text. Solving end-to-end Scene Text Recognition for Indian languages, often requires a pipeline of scene text detection, script identification followed by cropped word recognition. We use mean word recognition rate (WRR) and character recognition rate (CRR) over all the test images in the dataset. The WRR and CRR for an image are defined as follows:
@@
    where $S$, $D$, and $I$ with subscript $w$ and $c$ denote the number of substitutions, deletions, and insertions, respectively, at the word or character level. Further, $N_{w}^{\text{total}}$ and $N_{c}^{\text{total}}$ denote the total number of words and characters present in the ground-truth transcription in the image, respectively. It should be noted that due to false alarms (i.e., excessive insertions of words or characters in the recognized text) or in some rare cases of any missing annotation, the computed WRR or CRR values can become negative for few images. To keep the interpretation simple, we truncate such values to 0. Furthermore, standard WRR and CRR computations assume a fixed reading order; however, since reading order in scene text images can be subjective, we relax this constraint in our evaluation.
    Further, to make interpretation of results more clean, we also use standard precision, recall, and F1-score to evaluate end-to-end performance. Note that precision is defined as the number of correctly recognized words over the total number of recognized words, while recall is defined as the number of correctly recognized words over the total number of words in the ground truth.
Table 2: Summary of bounding box annotation in the BSTD, detailing the total, training, and test set distributions
@@
Table 3: Cropped word image data split for the BSTD, showing the number of annotated instances, train images, and test images across multiple languages. Here, Others correspond to those words which are not from the listed languages here (e.g., Meitei and Urdu). We plan to include their annotation in subsequent versions of our dataset. For scene-level language distribution statistics, please refer to Table 4.
@@
Figure 6: Distribution of word-level bounding box areas in the BSTD, plotted on a base-10 logarithmic scale.
Table 4: Scene-level language composition of the BSTD. The distribution highlights the multilingual nature of the dataset, with a significant majority of images containing English co-occurring with one or more Indian languages.
Language Composition #Images
English only 53
Indian languages only 1,958
English + one or more Indian languages 4,518
No annotatable scene text 53
Total 6,582